# Kinodynamic Task and Motion Planning using VLM-guided and Interleaved Sampling

### v1

@@ Abstract
Task and Motion Planning (TAMP) integrates high-level task planning with low-level motion feasibility, but existing methods are costly in long-horizon problems due to excessive motion sampling. While LLMs provide commonsense priors, they lack 3D spatial reasoning and cannot ensure geometric or dynamic feasibility. We propose a kinodynamic TAMP framework based on a hybrid state tree that uniformly represents symbolic and numeric states during planning, enabling task and motion decisions to be jointly decided. Kinodynamic constraints embedded in the TAMP problem are verified by an off-the-shelf motion planner and physics simulator, and a VLM guides exploring a TAMP solution and backtracks the search based on visual rendering of the states. Experiments on the simulated domains and in the real world show $32.14\%\sim 1166.67\%$ increased average success rates compared to traditional and LLM-based TAMP planners and reduced planning time on complex problems, with ablations further highlighting the benefits of VLM guidance.
@@ I INTRODUCTION
Robotic manipulation tasks, such as tabletop manipulations, require reasoning over both symbolic task decisions and continuous geometric feasibility. A robot must decide which action to perform—such as picking, placing, or stacking— and which object to grasp, which constitutes a discrete search process. Simultaneously, it must determine grasp poses, feasible end-effector configurations, and collision-free motion trajectories governed by continuous constraints. This class of problems is studied under the framework of Task and Motion Planning (i.e., TAMP), which combines high-level task planning with continuous action parameter binding and low-level motion planning [1, 2].
@@
When every high-level task plan has a feasible low-level motion solution, i.e., the downward refinement property, task plans can be computed independently of geometric constraints and later refined with continuous action parameters. Unfortunately, this assumption rarely holds in practice, where geometric infeasibility can invalidate a task plan. Hence, most TAMP systems interleave discrete and continuous reasoning. Early approaches follow two main paradigms: sequencing-first, which generates symbolic task plans first and then attempts to satisfy continuous constraints, or satisfaction-first, which samples feasible motions and integrates them into task planning [1]. The former suffers from repeated expensive constraint solving when samples are infeasible, while the latter wastes computation on many unused samples [3].
@@
With the emergence of large language models (LLMs), a new class of TAMP methods has been proposed. These methods leverage the commonsense knowledge and high-level reasoning abilities of pretrained models to generate problem formulation for TAMP [4], plausible task sequences [5], or even continuous parameters, such as placing positions, and motion failure reasoning [6]. However, due to the lack of their 3D spatial understanding, LLMs often struggle with precise motion reasoning and cannot reliably validate whether a plan is geometrically or physically feasible. Supplying them with high-dimensional numeric values like 6D poses or motion trajectories yields little benefit [5, 7]. Furthermore, existing TAMP approaches often ignore inertial or dynamic constraints in the planning environment [8]. This motivates the need to incorporate kinodynamic constraints into our framework, ensuring that generated motions are executable in the real world.
@@
To overcome these limitations, we propose a novel kinodynamic TAMP framework that interleaves sequencing-first and satisfaction-first approaches where task-level decisions and motion validations are coupled at every search-expansion step. This is abstracted through a hybrid search tree representation, where states and actions are grounded with sampled action parameters and immediately validated in simulation. An off-the-shelf motion planner and a physics simulator ensure kinodynamic feasibility, such as e.g., collisions, kinematics, grasp stability, or object stability. The VLM further evaluates resulting world states to bias exploration toward promising branches and enable backtracking to recover from unexpandable states. We validate our approach on two simulated and one real-world manipulation domains with varying problem complexity. Our method outperforms a traditional domain-independent TAMP planner [9] and an LLM-based TAMP planner [6] in terms of success rates and achieves shorter planning times for complex problems, in contrast to the baselines, which mostly fail. We also perform an ablation study to show that VLM guidance significantly improves the quality of backtracking decisions.
@@ II-A Task and Motion Planning (TAMP)
Traditional TAMP approaches are commonly categorized as sequencing-first, satisfaction-first, or interleaved. Sequencing-first methods first compute a symbolic task plan and then attempt to satisfy continuous constraints. Repeatedly solving constraint satisfaction problems after motion failures can be costly: these methods may suffer from planning-time explosion for long horizon tasks [9], require workspace discretization that limits scalability [10], or rely on manually specified constraints that are difficult to generalize across different problems and domains [3]. Satisfaction-first methods [11, 12] generate samples before symbolic search, effective when sampling is cheap. However, they suffer from combinatorial sample explosion, often producing many unnecessary samples.
Interleaved methods integrate sequencing and action parameter sampling within a unified search. For instance, sequencing-first planning can be combined with interleaved reasoning by pruning geometric samples making the task plan infeasible [13]. However, this is limited to refining a single fixed plan skeleton and does not explore diverse task plans. [14] constructs a reachability graph through backward search on a relaxed problem, which consists of useful actions to guide forward search. However, this graph approximates constraints independently and may therefore include inconsistent actions. More recent frameworks, such as [15], treat each plan skeleton as a top-level branch and employ MCTS for motion binding, but remain limited by random rollout policies that scale poorly in large domains. Our approach generates a discrete state graph to explore diverse task plans and expands the hybrid tree only with actions satisfying all constraints, while using a VLM for backtracking and a physics simulator for kinodynamic feasibility.
Learning-based approaches learn transition models [16], symbolic operators [17], or predicates [18] for TAMP. Recent works apply transformers [19] to predict the feasibility of plans or GNNs [20] to predict the relevance of sampled objects. However, data collection is costly and generalization is limited, whereas we leverage pre-trained VLMs with commonsense knowledge.
@@ II-B LLMs/VLMs for Task and Motion Planning
LLMs and VLMs have recently been applied to robotic planning, leveraging commonsense reasoning to sequence action primitives [21, 22, 23] or generate executable code that captures spatial reasoning and feedback loops [24]. They have also been combined with symbolic planning languages [25, 26, 27, 28, 29]. However, these methods are confined to task-level reasoning, without handling complex geometric or physical constraints.
LLMs and VLMs have been extended beyond simple task planning or skill sequencing to address TAMP problems. [30] employs LLMs to infer spatial relationships between objects, which are then passed to a TAMP planner. Other works generate problem formulations with LLMs/VLMs that are solved by external TAMP planners, with corrective re-prompting applied when failures occur [4, 31, 32, 33]. Similarly, [34] translates natural language constraints into Python APIs for an external TAMP planner, while [35] uses VLMs to infer open-world constraints—including both discrete action orders and continuous feasibility checks—through executable code integrated with an external planner. However, these approaches typically remain limited to coupling LLMs/VLMs with traditional TAMP planners.
[36] frames the search for valid LLM-generated programs and their continuous parameters as a constraint satisfaction problem. When violations occur, it enforces geometric and physical constraints through iterative re-prompting. LLMs can also directly produce task and motion plans: [5] generates batches of task plans refined with continuous parameters and expanded via MCTS, and [6] jointly generates symbolic sequences and continuous parameters, with re-prompting guided by motion planning feedback. However, unlike our approach, they rely solely on textual feedback for re-prompting and do not use visual cues to evaluate complex kinodynamic constraints during planning and replanning.
@@ II-C Kinodynamic Planning
Kinodynamic motion planning was first introduced by [8] to address both kinematic constraints, such as obstacle avoidance, and dynamic constraints, including limits on velocity, acceleration, force, and torque, to find time-optimal motion paths. [37] later proposed a randomized approach that extended RRTs to kinodynamic settings. Kinodynamic approaches have also been applied in TAMP. [38] integrated a sampling-based motion planner with a symbolic planner to jointly solve high-level task reasoning and low-level kinodynamic motion, thereby generating dynamically feasible trajectories. [39] employed differentiable physics engines to formulate planning as a gradient-based optimization problem, enabling reasoning over tool-use domains such as hitting and throwing. While classical kinodynamic planners focus on explicitly generating kinodynamic trajectories under robot dynamics, we instead focus on incorporating kinodynamic constraints into the TAMP problem.
@@ III PROBLEM FORMULATION
$P\equiv\langle\mathcal{O},\mathcal{H},\mathcal{A},\mathcal{T},h_{0},\mathcal{G}\rangle.$ (1)
Here, $\mathcal{O}$ is a finite set of environment objects, and $\mathcal{H}=\mathcal{S}\times\mathcal{X}$ denotes the hybrid state space. $\mathcal{S}$ is a finite collection of sets of symbolic states, where each symbolic state $s\in\mathcal{S}$ consists of PDDL [41] predicates such as (on bacon stove). $\mathcal{X}$ denotes the set of continuous world states (e.g., object poses, robot joint states, etc.). If $h=(s,x)\in\mathcal{H}$, then $x$ must be consistent with $s$, meaning that the continuous state $x$ satisfy the symbolic predicates in $s$. For instance, if $s=\{\texttt{(clear red)},\allowbreak\texttt{(on red green)},\allowbreak\texttt{(on-table green)},\allowbreak\texttt{(arm-empty)}\}$, then in $x$ the red block must not have any object on top, its pose must be located on the green block, the green block must be positioned on the table in a physically valid configuration, and no object should have contact with the arm gripper.
@@ IV-B Hybrid State Tree Expansion
The root node of the hybrid tree $T$, $h_{0}=(s_{0},x_{0})$, is defined by the initial PDDL state $s_{0}$ specified in the problem PDDL and an initial continuous state $x_{0}$. $x_{0}$ is instantiated by randomly sampling object states that satisfy the initial symbolic predicates $s_{0}$. Let the current hybrid node be $h_{t}=(s_{t},x_{t})$. Expanding from this node proceeds in three stages (the green block in Fig. 1):
@@ IV-B1 Candidate Action Generation
    Goal sampling: For pickup actions, a grasp pose is sampled using the object’s bounding box, offset by a sampled roll angle of the end-effector frame. For putdown actions, a feasible placement position within the target region and the roll angle of the end-effector frame are sampled randomly. Then collision detection is handled by an analytic narrow-phase solver for box–box collisions after an AABB-based sweep-and-prune broad phase by [45].
@@ IV-C2 VLM-Guided Backtracking
The constraint violation feedback is constructed from the outcomes of goal sampling, IK solver, motion planning, and candidate action simulation. Inspired by [6] and [36], we classify the feedback into four categories: (i) when the goal configuration collides with other objects, (ii) when no feasible IK solution exists, (iii) when the motion planner fails to produce a collision-free trajectory, and (iv) when a trajectory is found but grasp or release constraints are violated during execution. This structured feedback, together with the images of the current node $h_{t}$, helps the VLM identify the cause of failure using both textual feedback and visual cues. Based on this information, the VLM then selects an appropriate node to backtrack to within the JSON tree, say $h_{r}=(s_{r},x_{r})$. From $h_{r}$, step IV-B1, step IV-B2, and step IV-B3 in Sec. IV-B are repeated until either a goal state $h_{g}\in\mathcal{G}$ is reached or the timeout is exceeded.
By combining randomized retries with VLM-guided backtracking, our framework efficiently recovers from unexpandable nodes and explores promising alternative plan skeletons in the large hybrid search space. For example, suppose repeated attempts in the same region are likely to fail, such as in a cluttered workspace or blocked access. In that case, the VLM can backtrack to a node that enables clearing or repositioning other objects first.
@@ V-A Experimental Setup
All task and motion planning experiments were conducted on a workstation with a 13th Gen Intel Core i9-13900K CPU and an NVIDIA GeForce RTX 4090 GPU. We employed GPT-4o  [^1] https://openai.com from OpenAI as the VLM, with temperature fixed to 0.0. For generating symbolic plan skeletons, we used a top-$k$ symbolic planner [42] based on Fast-Downward [44] combined with the $K^{\ast}$ algorithm [43], with the number of plans fixed to $k=30$. Genesis [45] was used as the physics simulator and photo-realistic renderer. A planning timeout of 600 seconds was imposed for all experiments, with planning proceeding until a goal was found or the timeout was reached. After the planning was terminated, the resulting simulator state was checked against the PDDL goal conditions; if satisfied, the trial was marked as a success; otherwise, a failure.
@@
  • Kitchen: This domain is a modified version of the Kitchen domain used in PDDLStream [9]. As in Fig. 3, it contains six colored food objects (radish, egg, bacon, chicken, celery, apple), a sink (red cuboid), and a stove (blue cuboid). A food block becomes Cooked only after being cleaned on the sink and then cooked on the stove. The goal is to make a randomly chosen $n$ food objects into Cooked state. To increase problem difficulty, we added gray distractor blocks around the initial positions of the food items.
@@ V-B Performance Analysis
For both domains, we varied the number of target objects $3\leq n\leq 6$. For each value of $n$, we randomly generated 10 problem instances and measured both success rate and planning time. Table. I shows the results for each domain. We quantitatively compared four approaches:
@@
    Ours w/o VLM Backtrack: Our method without VLM guidance for replanning. When a failure occurs, we return to a simple BFS-based strategy that selects the first unvisited discrete node and resumes search from the corresponding hybrid node.
@@
    PDDLStream[9]: A domain-independent TAMP baseline that integrates symbolic planning with black-box sampling functions (e.g., grasp sampler, IK solver, RRT-Connect motion planner, etc.). We compare against its ”adaptive” algorithm. Search-sample ratio is fixed to 1.
@@
    LLM3[6]: An LLM-based TAMP framework where the LLM directly predicts symbolic action sequences and continuous parameters. We compare against the variant ”LLM3 Backtrack”, where the LLM selects a previous action for backtracking and replans from that point upon failure.
@@
Comparisons: Our method outperforms baselines, achieving an average success rate of 92.5% in Blocksworld and 95% in Kitchen. In contrast, PDDLStream reports only 45% and 7.5% success rates in Blocksworld and Kitchen, respectively. Moreover, its planning time grows exponentially as $n$ increases, eventually leading to timeouts at $n=6$ in Blocksworld, $n=4,5,6$ in Kitchen for all ten problem instances. PDDLStream employs streams, which are black-box sampling functions that generate continuous parameters and certify constraints such as no collisions. The outputs of streams, called ”stream objects”, are added to the symbolic state and reused when searching for new plans. As more stream objects accumulate, the planning state space grows significantly, which slows down the process and often leads to inefficiency [3].
@@
LLM3 achieved average success rates of 70% in Blocksworld and 25% in Kitchen, but suffered timeouts at $n=4,6$ in Kitchen for all problem instances and at $n=5$ for 90% of problem instances. This is mainly due to collisions when placing food blocks on the stove, as LLMs struggle to predict collision-free poses and continue to produce incorrect outputs even after replanning. In Blocksworld, LLM3’s success rate rapidly decreased at $n=6$ because, as the number of blocks increased, the model more frequently predicted misordered stacking sequences, resulting in an average of 42 backtrack trials at $n=6$.
@@
Ablation Study: We conducted an ablation study to evaluate the effectiveness of VLM-guided backtracking. As shown in Table I, our method with VLM-guided backtracking achieved 23.33% and 8.57% increased average success rates in Blocksworld and Kitchen, respectively, compared to the version without VLM guidance. The planning time showed no significant difference between the two settings. This is because VLMs leverage commonsense knowledge and visual reasoning to identify the cause of failure in the current state and backtrack to a feasible node within one or two attempts. In contrast, without VLM-guided backtracking, the planner often repeatedly attempted backtracking until reaching a timeout.
@@
Large Task Space vs. Large Motion Space: Blocksworld exemplifies a large task space, as its complexity arises from the high branching factor induced by arbitrary stacking relations and discrete actions such as stack and unstack [20, 48]. By contrast, the Kitchen domain is characterized by a large motion space, where tight manipulation constraints make it challenging to grasp food objects in clutter due to distractors and to place multiple items on the stove or sink without collisions [15, 48].
@@
Our ablation also confirms this distinction: VLM-guided backtracking yields a larger performance gain in Blocksworld than in Kitchen. In Blocksworld, where difficulty is primarily due to symbolic branching, VLM backtracking effectively explores alternative discrete states and mitigates task-level search complexity. In the Kitchen, however, the primary challenge lies in continuous motion feasibility under cluttered initial positions and constrained placements, so backtracking to a different discrete action has less impact. This demonstrates that the effectiveness of our VLM guidance is closely tied to whether domain complexity arises from task-space branching or motion-space feasibility.

### v2

@@ Abstract
Task and Motion Planning (TAMP) integrates high-level task planning with low-level motion feasibility, but existing methods are costly in long-horizon problems due to excessive motion sampling. While LLMs provide commonsense priors, they lack 3D spatial reasoning and cannot ensure geometric or dynamic feasibility. We propose a kinodynamic TAMP planner based on a hybrid state tree that uniformly represents symbolic and numeric states during planning, enabling task and motion decisions to be jointly decided. Kinodynamic constraints embedded in the TAMP problem are verified by an off-the-shelf motion planner and physics simulator, and a VLM guides exploring a TAMP solution and backtracks the search based on visual rendering of the states. Experiments on the simulated domains and in the real world show $32.14\%\sim 1166.67\%$ increased average success rates compared to traditional and LLM-based TAMP planners and reduced planning time on complex problems, with ablations further highlighting the benefits of VLM backtracking. More details are available at https://graphics.ewha.ac.kr/kinodynamicTAMP/.
@@ I INTRODUCTION
Robotic manipulation tasks, such as tabletop manipulations, require reasoning over both symbolic task decisions and continuous geometric feasibility. A robot must decide which action to perform—such as picking or placing— and which object to grasp, which constitutes a discrete search process. Simultaneously, it must determine grasp poses, feasible end-effector configurations, and collision-free motion trajectories governed by continuous constraints. This class of problems is studied under the framework of Task and Motion Planning (i.e., TAMP), which combines high-level task planning with continuous action parameter binding and low-level motion planning [1].
@@
When every high-level task plan has a feasible low-level motion solution, i.e., the downward refinement property, task plans can be computed independently of geometric constraints and later refined with continuous action parameters. Unfortunately, this assumption rarely holds in practice, where geometric infeasibility can invalidate a task plan. Hence, most TAMP systems interleave discrete and continuous reasoning. Early approaches follow two main paradigms: sequencing-first, which generates symbolic task plans first and then attempts to satisfy continuous constraints, or satisfaction-first, which samples feasible motions first and integrates them into task planning [1]. The former suffers from repeated expensive constraint solving when samples are infeasible, while the latter wastes computation on many unused samples [2].
@@
With the emergence of large language models (LLMs), a new class of TAMP solvers has been proposed. These methods leverage the commonsense knowledge and high-level reasoning abilities of pretrained models to generate problem formulations for TAMP [3], plausible task sequences [4], and even continuous parameters, such as placement positions and motion failure reasoning [5]. However, due to the lack of their 3D spatial understanding, LLMs often struggle with precise motion reasoning and cannot reliably validate whether a plan is geometrically or physically feasible. Supplying them with high-dimensional numeric values like 6D poses or motion trajectories yields little benefit [4, 6]. Furthermore, existing TAMP approaches often ignore inertial or dynamic constraints in the planning environment. This motivates the need to incorporate kinodynamic constraints into our framework, ensuring that generated motions are executable in the real world.
@@
To overcome these limitations, we propose a novel kinodynamic TAMP framework that interleaves a sequencing-first and a satisfaction-first approach, coupling task-level decisions and motion validations at every search-expansion step. This is abstracted through a hybrid search tree representation, where states and actions are grounded with sampled action parameters and immediately validated in physics simulation. An off-the-shelf motion planner and a physics simulator ensure kinodynamic feasibility, such as e.g., collisions, kinematics, grasp stability, and object stability. The VLM further evaluates resulting world states to bias exploration toward promising branches and enable backtracking to recover from unexpandable states. We validate our approach on two simulated and one real-world manipulation domains with varying problem complexity. Our method outperforms a traditional domain-independent TAMP planner [7] and an LLM-based TAMP planner [5] in terms of success rates and achieves shorter planning times on complex problems, whereas the baselines mostly fail. We also perform an ablation study to show that VLM guidance significantly improves the quality of backtracking decisions.
@@ II-A Task and Motion Planning (TAMP)
Traditional TAMP approaches are commonly categorized as sequencing-first, satisfaction-first, or interleaved. Sequencing-first methods first compute a symbolic task plan and then attempt to satisfy continuous constraints. Repeatedly solving constraint satisfaction problems after motion failures can be costly: these methods may suffer from planning-time explosion for long horizon tasks [7], require workspace discretization that limits scalability [8], or rely on manually specified constraints that are difficult to generalize across different problems and domains [2]. Satisfaction-first methods [9, 10] generate samples before symbolic search, effective when sampling is cheap. That said, they suffer from combinatorial sample explosion, often producing many unnecessary samples.
Interleaved methods integrate sequencing and action parameter sampling within a unified search. For instance, sequencing-first planning can be combined with interleaved reasoning by pruning geometric samples that make the task plan infeasible [11]. However, this is limited to refining a single fixed plan skeleton and does not explore diverse task plans. More recent frameworks, such as [12], treat each plan skeleton as a top-level branch and employ MCTS for motion binding, but remain limited by random rollout policies that scale poorly in large domains. Our approach generates a discrete state graph to explore diverse task plans and expands the hybrid tree only with actions that satisfy all constraints, validated by a physics simulator for kinodynamic feasibility.
Learning-based approaches learn symbolic operators [13], or predicates [14] for TAMP. Recent works apply transformers [15] to predict the feasibility of plans or GNNs [16] to predict the relevance of sampled objects. However, data collection is costly, and generalization is limited, whereas we leverage pre-trained VLMs with commonsense knowledge.
@@ II-B LLMs/VLMs for Task and Motion Planning
LLMs and VLMs have recently been applied to robotic planning, leveraging commonsense reasoning to sequence high-level action primitives [17, 18, 19] or generate robot executable code that captures spatial reasoning and feedback loops [20]. They have also been combined with symbolic planning languages [21, 22, 23, 24]. However, these methods are restricted to task-level reasoning, without handling complex geometric or physical constraints.
LLMs and VLMs have been extended beyond simple task planning or skill sequencing to address TAMP problems. [25] employs LLMs to infer spatial relationships between objects, which are then passed to a TAMP planner. Other works generate problem formulations with LLMs/VLMs that are solved by external TAMP planners, with corrective re-prompting applied when failures occur [3, 26, 27, 28]. Similarly, [29] translates natural language constraints into Python APIs for an external TAMP planner, while [30] uses VLMs to infer open-world constraints—including both high-level action orders and continuous feasibility checks—through executable code integrated with an external planner. However, these approaches typically remain limited to coupling LLMs/VLMs with traditional TAMP planners.
[31] frames the search for valid LLM-generated Python programs and their continuous parameters as a constraint satisfaction problem. When violations occur, it enforces geometric and physical constraints through iterative re-prompting. LLMs can also directly produce task and motion plans: [4] generates batches of task plans refined with continuous parameters and expanded via MCTS, and [5] jointly generates symbolic sequences and continuous parameters, with re-prompting guided by motion planning feedback. However, unlike our approach, they rely solely on textual feedback for re-prompting and do not use visual cues to evaluate complex kinodynamic constraints during planning and replanning.
@@ II-C Kinodynamic Planning
Kinodynamic motion planning was first introduced by [32] to address both kinematic constraints, such as obstacle avoidance, and dynamic constraints, including limits on velocity, acceleration, force, and torque, to find time-optimal motion paths. [33] later proposed a randomized approach that extended RRTs to kinodynamic settings. Kinodynamic approaches have also been applied in TAMP. [34] integrated a sampling-based motion planner with a symbolic planner to jointly solve high-level task reasoning and low-level kinodynamic motion, thereby generating dynamically feasible trajectories. [35] employed differentiable physics engines to formulate planning as a gradient-based optimization problem, enabling reasoning over tool-use domains such as hitting and throwing. While classical kinodynamic planners focus on explicitly generating kinodynamic trajectories under robot dynamics, we instead focus on incorporating kinodynamic constraints and physical stability into the TAMP problem.
@@ III PROBLEM FORMULATION
$P\equiv\langle\mathcal{H},\mathcal{O},\mathcal{A},\mathcal{T},h_{0},\mathcal{G}\rangle.$ (1)
Here, $\mathcal{H}=\mathcal{S}\times\mathcal{X}$ denotes the hybrid state space, and $\mathcal{O}$ is a finite set of environment objects. $\mathcal{S}$ is a finite set of symbolic states, where each symbolic state $s\in\mathcal{S}$ consists of PDDL [37] predicates such as (on bacon stove). $\mathcal{X}$ denotes the set of continuous world states (e.g., object poses, robot joint states, etc.). If $h=(s,x)\in\mathcal{H}$, then $x$ must be consistent with $s$, meaning that the continuous state $x$ satisfy the symbolic predicates in $s$. For instance, if $s=\{\texttt{(clear red)},\allowbreak\texttt{(on red green)},\allowbreak\texttt{(on-table green)},\allowbreak\texttt{(arm-empty)}\}$, then in $x$, the red block must not have any object on top, its position must be located on top of the green block, the green block must be positioned on the table in a physically valid configuration, and no object should have contact with the arm gripper.
@@ IV-B Hybrid State Tree Expansion
The root node of the hybrid state tree $T$, $h_{0}=(s_{0},x_{0})$, is defined by the initial PDDL state $s_{0}$ specified in the problem PDDL and an initial continuous state $x_{0}$. $x_{0}$ is instantiated by randomly sampling object states that satisfy the initial symbolic predicates $s_{0}$. Let the current hybrid node be $h_{t}=(s_{t},x_{t})$. Expanding from this node proceeds in three stages (the green block in Fig. 1):
@@ IV-B1 Candidate Action Generation
    Goal sampling: For pickup actions, a grasp pose is sampled based on the object’s bounding box, with a randomly sampled roll angle applied to the end-effector frame. For putdown actions, a feasible placement position within the target region and the roll angle of the end-effector frame are sampled randomly. Then collision detection is handled by an analytic narrow-phase solver for box–box collisions after an AABB-based sweep-and-prune broad phase by [41].
@@ IV-C2 VLM-Guided Backtracking
The constraint violation feedback is constructed from the outcomes of goal sampling, IK solver, motion planning, and candidate action simulation. Inspired by [5] and [31], we classify the feedback into four categories: (i) when no feasible IK solution exists, (ii) when the goal configuration collides with other objects, (iii) when the motion planner fails to produce a collision-free trajectory, and (iv) when a trajectory is found but grasp or release constraints are violated during execution. This structured feedback, together with the images of the current node $h_{t}$, helps the VLM identify the cause of failure using both textual feedback and visual cues. Based on this information, the VLM then selects an appropriate node to backtrack to within the JSON tree, say $h_{r}=(s_{r},x_{r})$. From $h_{r}$, step IV-B1, step IV-B2, and step IV-B3 in Sec. IV-B are repeated until either a goal state $h_{g}\in\mathcal{G}$ is reached or the timeout is exceeded.
By combining randomized retries with VLM-guided backtracking, our framework efficiently recovers from unexpandable nodes and explores promising alternative states in the large hybrid search space. For example, suppose repeated attempts in the same region are likely to fail, such as in a cluttered workspace or blocked access. In that case, the VLM can backtrack to a node that enables clearing or repositioning other objects first.
@@ V-A Experimental Setup
All TAMP experiments were conducted on a workstation with a 13th Gen Intel Core i9-13900K CPU and an NVIDIA GeForce RTX 4090 GPU. We employed GPT-4o  [^1] https://openai.com from OpenAI as the VLM, with temperature fixed to 0.0. For generating symbolic plan skeletons, we used a top-$k$ symbolic planner [38], with the number of plans fixed to $k=30$. Genesis [41] was used as the physics simulator and photo-realistic renderer. A planning timeout of 600 seconds was imposed for all experiments, with planning proceeding until a goal was found or the timeout was reached. After the planning was terminated, the resulting simulator state was checked against the PDDL goal conditions; if satisfied, the trial was marked as a success; otherwise, a failure.
@@
  • Kitchen: This domain is a modified version of the Kitchen domain used in PDDLStream [7]. As in Fig. 3, it contains 6 colored food objects (radish, egg, bacon, chicken, celery, apple), a sink (red cuboid), and a stove (blue cuboid). A food block becomes Cooked only after being cleaned on the sink and then cooked on the stove. The goal is to make a randomly chosen $n$ food objects into Cooked state. To increase problem difficulty, we added gray distractor blocks around the initial positions of the food objects.
@@ V-B Performance Analysis
For both domains, we varied $3\leq n\leq 6$. For each value of $n$, we randomly generated 10 problem instances and measured both success rate and planning time. Table. I shows the results for each domain. We quantitatively compared 4 approaches:
@@
    Ours w/o VLM Backtrack: Our method without VLM guidance for backtracking. When a failure occurs, we return to a simple BFS-based strategy that selects the first unvisited discrete node and resumes search from the corresponding hybrid node.
@@
    PDDLStream[7]: A domain-independent traditional TAMP baseline that integrates symbolic planning with black-box sampling functions (e.g., grasp sampler, IK solver, RRT-Connect motion planner, etc.). We compare against its ”adaptive” algorithm. Search-sample ratio is fixed to 1.
@@
    LLM3[5]: An LLM-based TAMP baseline where the LLM directly predicts symbolic action sequences and continuous parameters. We compare against the variant ”LLM3 Backtrack”, where the LLM selects a previous action for backtracking and replans from that point upon failure.
@@
Comparisons: Our method outperforms baselines, achieving an average success rate of 92.5% in Blocksworld and 95% in Kitchen. In contrast, PDDLStream reports success rates of only 45% and 7.5% in Blocksworld and Kitchen, respectively. Moreover, its planning time grows exponentially as $n$ increases, eventually leading to timeouts at $n=6$ in Blocksworld, $n=4,5,6$ in Kitchen for all ten problem instances. PDDLStream employs streams, which are black-box sampling functions that generate continuous parameters and certify constraints, such as collision-free motion. The outputs of streams, called ”stream objects”, are added to the symbolic state and reused when searching for new plans. As more stream objects accumulate, the planning state space grows significantly, slowing the process and often leading to inefficiency [2].
@@
LLM3 achieved average success rates of 70% in Blocksworld and 25% in Kitchen, but suffered timeouts at $n=4,6$ in Kitchen for all problem instances and at $n=5$ for 90% of problem instances. In Blocksworld, the success rate drops at $n=6$, where the LLM increasingly predicts misordered stacking sequences, resulting in an average of 42 backtrack trials. In Kitchen, failures arise from collisions when placing food blocks on the stove, as LLMs struggle to predict collision-free poses and often continue producing infeasible continuous action parameters outputs even after replanning.
The comparison with the baselines clarifies the role of each component in our pipeline. The discrete state graph generated by the top-$k$ planner explicitly bounds the symbolic search space, preventing the planning-time explosion caused by uncontrolled search space growth observed in PDDLStream. Moreover, hybrid state tree expansion with motion planning and physics-based rollout grounds each action in geometric and dynamic feasibility, mitigating the geometric reasoning limitations of purely LLM-based TAMP planners.
@@
Ablation Study of VLM-guided Backtracking: We conducted an ablation study to evaluate the effectiveness of VLM-guided backtracking. As shown in Table I, our method with VLM-guided backtracking achieved 23.33% and 8.57% increases in average success rates in Blocksworld and Kitchen, respectively, compared to the version without VLM guidance. This is because VLMs leverage commonsense knowledge and visual reasoning to identify the cause of failure in the current state and backtrack to a feasible node within one or two attempts. In contrast, without VLM-guided backtracking, the planner often repeatedly attempted backtracking until reaching a timeout. While incorporating VLM-guided backtracking slightly increased planning time due to additional VLM calls, the overall difference was marginal.
@@
Large Task Space vs. Large Motion Space: Blocksworld exemplifies a large task space, as its complexity arises from the high branching factor induced by arbitrary block stacking configurations and discrete actions such as stack and unstack [16, 44]. By contrast, the Kitchen domain is characterized by a large motion space, where tight manipulation constraints make it challenging to grasp food objects in cluttered environments with distractors and to place multiple items on the stove or sink without collisions [12, 44].
@@
Our ablation also confirms this distinction: VLM-guided backtracking yields a larger performance gain in Blocksworld than in Kitchen. In Blocksworld, where difficulty is primarily due to symbolic branching, VLM backtracking effectively explores alternative discrete states and mitigates task-level search complexity. In the Kitchen, however, the primary challenge lies in continuous motion feasibility under cluttered initial positions and constrained placements, so backtracking to a different discrete action has less impact. This demonstrates that VLM-guided backtracking is particularly effective in domains characterized by large task spaces, where symbolic branching dominates overall complexity.